OpenStudy (anonymous):

Find two consecutive odd integers such that the square of the first, added to 3 times the second is 24.

OpenStudy (anonymous):

@david111

OpenStudy (anonymous):

Odd numbers: 1,3,5,7,9,11,13,15

OpenStudy (anonymous):

let's assume that those 2 consecutive odd integers are \(m\) and \(m+2\) so you will come up with this equation:\[m^2+3(m+2)=24\]this is a quadratic :)

OpenStudy (anonymous):

3 and 5

OpenStudy (anonymous):

Lets call our integers i and j. Any 2 consecutive odd integers will always be 2 apart, so we can first say that: \[j=i+2\] From the remainder of the question, we can say that: \[i ^{2}+3j=24\] Substituting the first equation in the second gives: \[i ^{2}+3*(i+2)=24\] Rearranging and expanding brackets gives us a quadratic equation: \[i ^{2}+3i-18=0\] Factorising it gives: (i-3)*(i+6)=0 Therefore i is either 3 or -6. Since i is odd according to the original question, i must be 3. Since j is the following odd integer after i, j must be 5. Result: The 2 integers are 3 and 5.

OpenStudy (anonymous):

thats what i said

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome